LM equation specified with Money demand L = 0.20 Y – 4i and money supply is $144.

IS equations for a two sector model specified as C= 90£ + 0.75 Y and I= 140£ – 5i and the second with C= 90£ + 0.75Y and I= 120£ – 3i.

 

    What does LM and IS means? Calculate the LM and the two IS equations.

    Plot LM and the two IS curves

    Plot the new LM equation and label it LM1 for the 15£ increase in Ms.

     Use the IS equations and find output when the Ms increases to 159£. For which IS equation is there a greater decrease in the rate of interest?

     Find investment and consumption spending when Ms is 144£ and when Ms is 159£.

     The increase in the Ms causes a shift in the LM or the IS curve? Why? And what is called?

   What happened to IS and LM curves is G increased by 10£?

 


1
Expert's answer
2021-04-19T18:49:26-0400

The IS-LM model, which stands for "investment-savings" (IS) and "liquidity preference-money supply" (LM) is a Keynesian macroeconomic model that shows how the market for economic goods (IS) interacts with the loanable funds market (LM) or money market.

LM is:

0.20 Y – 4i = 144,

Y = 720 + 20i.

IS1 is:

Y = 90 + 0.75 Y + 140 – 5i,

Y = 920 - 20i

IS2 is:

Y = 90 + 0.75Y + 120 – 3i,

Y = 840 - 12i.

LM curve is upward-sloping, and IS curves are downward-sloping.

If Ms increases by 15, then LM2 is:

0.20 Y – 4i = 159,

Y = 795 + 20i.

For IS1:

795 + 20i = 920 - 20i,

40i = 125,

i = 3.125%.

Y = 920 - 20×3.125 = £857.5.

For IS2:

795 + 20i = 840 - 12i,

32i = 45,

i = 1.41%.

Y = 840 - 12×1.41 = £823.08.

At higher Ms investment and consumption spending are higher.

The increase in the Ms causes a shift in the LM curve.

If G increased by 10£, then IS increased.